mr. jiang is building a deck on his house. he can calculate the perimeter of the desk using this formula : P=4s
solve the equation to show the length of a side-that is, solve for s.
is mr. jiangs deck in the shape of a square or rectangle? explain your answer

Answer by ewatrrr(24785)


Hi
mr. jiang is building a deck on his house.
he can calculate the perimeter of the desk using this formula : P=4s
P = 4s then s = P/4
is mr. jiangs deck in the shape of a square or rectangle? explain your answer
In general for rectangle P = 2*L + 2*W
If the length is equal to the width... then P = 4*s
Mr Jiang's deck is a square deck